OpenStudy (lisa100):

what is the solution set of the equation log2x+log2(x-3)=2

OpenStudy (reemii):

\(\log (2x) + \log(2(x-3)) = 2\) like this?

OpenStudy (campbell_st):

is the question using \[\log_{2}(x) + \log_{2}(x -3) = 2\]

OpenStudy (lisa100):

it is kind of like log (subscript) 2

OpenStudy (anonymous):

I think campbell_st has it right.

OpenStudy (lisa100):

yes campbell_st

OpenStudy (lisa100):

how do i solve it?

OpenStudy (campbell_st):

ok... so simplify the logs 1st adding logs means they can be written as 1 and mean multiplying so you have \[\log_{2}(x \times (x-3)) = 2\] or \[\log_{2}(x^2 - 3x) = 2\] now raise each part as a power of 2... to remove the logs \[2^{\log_{2}(x^2 - 3x)} = 2^2...or..... x^2 - 3x = 4\] so you need to solve the quadratic \[x^2 - 3x - 4 = 0\]

OpenStudy (campbell_st):

hope it helps

OpenStudy (lisa100):

thanks it does!

OpenStudy (lisa100):

so is the solution set (-1,4)

OpenStudy (campbell_st):

well they are the solutions to the quadratic... you can check by substituting into the original equation... \[\log_{2}(x^2 - 3x)\] when x = -1 you get \[\log_{2}(1 - -3) = \log_{2}(4) = \log_{2}(2^2) = 2\] so that works and x = 4 \[\log_{2}(4^2 - 3 \times 4) = \log_{2}(4) .... \] and you can see the substitution is going the same way as -1 ... so I'd say it works....